OpenStudy (anonymous):

Which choice be the equation of a line parallel to the line represented by the equation? A. Y=-1/2×+4 B.y=2×+7 C.y=-2×+1 D.y=3/2×-5

OpenStudy (anonymous):

The is equation for y=2×-3

OpenStudy (anonymous):

Which choice be the equation of a line parallel to the line represented by the equation? Y=2×-3 A. Y=-1/2×+4 B.y=2×+7 C.y=-2×+1 D.y=3/2×-5

OpenStudy (steve816):

Parallel lines have the same slope. And since the slope is 2x, the answer is B

OpenStudy (steve816):

\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Huge\bf\color{#FF0000}W\color{#FF4900}e\color{#FF9200}l\color{#FFDB00}c\color{#FFff00}o\color{#B6ff00}m\color{#6Dff00}e\color{#24ff00}~\color{#00ff00}t\color{#00ff49}o\color{#00ff92}~\color{#00ffDB}O\color{#00ffff}p\color{#00DBff}e\color{#0092ff}n\color{#0049ff}S\color{#0000ff}t\color{#2400ff}u\color{#6D00ff}d\color{#B600ff}y\color{#FF00ff}!\)\(~~~~~~~~~~~~~~~~~~~~~~~~~~\Huge{\color{red}{\heartsuit}\color{orange}{\bigstar}\color{yellow}{\heartsuit}\color{green}{\bigstar}\color{blue}{\heartsuit}\color{purple}{\bigstar}\color{hotpink}{\heartsuit}}\)\(\Large\bf\color{Lime}{Please,\ don't\ hesitate\ to\ ask\ questions!}\)\(\Large\bf\color{magenta}{We're\ here\ to\ help!}\Large\color{magenta}{\ddot\smile}\)

OpenStudy (anonymous):

Can you please help with 6 more questions